# ChatThero: An LLM-Supported Chatbot for Behavior Change and Therapeutic Support in Addiction Recovery

### v1

@@ Abstract
Substance use disorders (SUDs) affect over 36 million people worldwide, yet few receive effective care due to stigma, motivational barriers, and limited personalized support. Although large language models (LLMs) show promise for mental-health assistance, most systems lack tight integration with clinically validated strategies, reducing effectiveness in addiction recovery. We present ChatThero, a multi-agent conversational framework that couples dynamic patient modeling with context-sensitive therapeutic dialogue and adaptive persuasive strategies grounded in cognitive behavioral therapy (CBT) and motivational interviewing (MI). We build a high-fidelity synthetic benchmark spanning Easy, Medium, and Hard resistance levels, and train ChatThero with a two-stage pipeline comprising supervised fine-tuning (SFT) followed by direct preference optimization (DPO). In evaluation, ChatThero yields a 41.5% average gain in patient motivation, a 0.49% increase in treatment confidence, and resolves hard cases with 26% fewer turns than GPT-4o, and both automated and human clinical assessments rate it higher in empathy, responsiveness, and behavioral realism. The framework supports rigorous, privacy-preserving study of therapeutic conversation and provides a robust, replicable basis for research and clinical translation.  [^1] Upon acceptance, we will release our training and evaluation code as well as data construction scripts. Due to the sensitive and potentially identifiable nature of addiction narratives, the full dataset and synthetic profiles will not be released.
ChatThero: An LLM-Supported Chatbot for Behavior Change and Therapeutic Support in Addiction Recovery
Junda Wang thanks: indicates equal contribution 1,2, Zonghai Yao11footnotemark: 1 1,2, Zhichao Yang2, Lingxi Li2, Junhui Qian3, Hong Yu1,2,3,4 1Center for Healthcare Organization and Implementation Research, VA Bedford Health Care 2Manning College of Information and Computer Sciences, University of Massachusetts Amherst 3Miner School of Computer and Information Sciences, University of Massachusetts Lowell 4Department of Medicine, University of Massachusetts Medical School {jundawang, zonghaiyao}@umass.edu
@@ 1 Introduction
Figure 1: We evaluate several LLM-based chatbots on their ability to express motivation and confidence in simulated addiction recovery dialogues. Each point represents the average score (1–5 scale). Our ChatThero (green; built on Qwen-7B), consistently outperforms all baselines across patient difficulty levels. The performance gap is especially large for medium and hard patients—groups where other models show low confidence and motivation, while ChatThero remains notably stronger, demonstrating greater resilience and communicative effectiveness in challenging scenarios.
Drug addiction represents a critical global health crisis affecting millions worldwide, driving compulsive substance-seeking behaviors despite severe medical, psychological, and social consequences. In the United States alone, approximately 2.5 million individuals currently suffer from opioid use disorders (OUD), with over 66 million individuals reporting recent illicit drug use (Recovery Team, 2023). Substance use disorders (SUDs) remain notoriously challenging to manage due to their chronic, relapsing nature; approximately 80–90% of individuals with opioid addiction relapse within the first year without adequate, sustained support (Wooten and Pieper, ). Particularly vulnerable are individuals with co-occurring mental health conditions, who face heightened risks of relapse, overdose, homelessness, and criminal justice involvement (Jones and McCance-Katz, 2019; Di Lorenzo et al., 2014). Despite the existence of effective treatments such as medication-assisted therapy (MAT) and cognitive behavioral therapy (CBT), fewer than a quarter of individuals with OUD receive any form of professional intervention (Cloyes et al., 2010). Significant barriers, including pervasive social stigma, financial limitations, and geographical isolation from specialized care, severely restrict access and undermine patient engagement (Sinha, 2011; Cloyes et al., 2010).
While recent advancements in Large Language Models (LLMs) Achiam et al. (2023); Jaech et al. (2024); Grattafiori et al. (2024); Anthropic (2024); Team et al. (2023); Yang et al. (2025); Guo et al. (2025) have revolutionized multiple sectors through human-level conversational fluency and sophisticated decision support Zhao et al. (2023); Xi et al. (2025); Wang et al. (2024); Li et al. (2024b); Guo et al. (2024); Tran et al. (2025), the AI research community has largely under-addressed the urgent, complex needs associated with substance use treatment and recovery. Traditional AI-based health interventions often lack mechanisms for modeling critical psychological dynamics such as real-time motivational fluctuations, sustained therapeutic engagement, and the personalized, adaptive support necessary for individuals battling addiction. To date, AI-driven solutions specifically tailored to the long-term behavioral changes required in addiction recovery remain scarce Yao and Yu (2025); Laymouna et al. (2024); Wang et al. (2025).
To address these critical challenges, we propose ChatThero, a novel AI-driven conversational framework explicitly designed for addiction recovery. Distinct from general-purpose conversational agents, ChatThero uniquely integrates multi-agent simulations built on dynamically evolving patient profiles and adaptive memory structures (Figure 2, left). Patient profiles—synthesized from anonymized real-world narratives—capture nuanced psychological traits (e.g., impulsivity, distrust), detailed substance use histories, and personalized treatment barriers (e.g., stigma, emotional reliance), enabling contextually rich and realistic patient modeling. Our framework utilizes a two-stage training pipeline: initially employing supervised fine-tuning (SFT) with synthetic therapeutic dialogues grounded in CBT and motivational interviewing (MI), followed by Direct Preference Optimization (DPO) guided by expert and AI feedback to dynamically refine persuasive strategies (Figure 2, middle and right). Evaluations across simulated dialogues demonstrate substantial effectiveness, with ChatThero significantly surpassing advanced baselines (GPT-4o, LLaMA3-8B, and Qwen variants) particularly in challenging scenarios: it improves patient motivation scores from 2.39 to 4.10 and confidence from 1.52 to 3.19 in hard recovery cases (Figure 1). Human evaluators further confirm ChatThero’s superior empathetic responsiveness (4.93/5) and clinical relevance (4.61/5), highlighting its potential for sustained therapeutic engagement in realistic, resistance-heavy addiction recovery interactions.
Our work makes the following novel contributions:
  • Introducing the first dedicated computational framework and simulation benchmark specifically for addiction treatment, thereby filling a significant void in AI-driven social impact research.
  • Proposing a novel therapeutic conversational framework that integrates dynamic profiles, long-term adaptive memory, and multi-agent interactions grounded in clinically validated strategies (CBT, MI), specifically tailored for sustained behavior change in high-relapse contexts.
  • Adhering to rigorous ethical standards by employing fully anonymized, synthesized patient profiles, and simulation-based evaluations to ensure safe, ethical exploration of sensitive therapeutic interactions without compromising individual privacy.
  • Our documented methods, benchmark construction scripts, and synthetic dataset generation tools will be openly available to facilitate reproducibility and further research. The flexible and extensible architecture of ChatThero can be readily adapted for other social impact domains involving long-term behavioral change.
@@ 2 Related Works
2.1 Substance Use Disorders: Treatment Gaps and Behavioral Challenges
Evidence-based treatments for substance use disorders, including MAT and CBT, have demonstrated significant efficacy in clinical settings by reducing relapse rates and improving psychological functioning (SAM, 2023; Wooten and Pieper, ; McHugh et al., 2010). However, translating these successes into real-world practice remains challenging. Only 22.1% of individuals with past-year opioid use disorder receive treatment, and fewer than half remain engaged after six months (Cloyes et al., 2010; SAM, 2023). Common barriers include social stigma, financial constraints, geographic limitations, and patient ambivalence (Sinha, 2011; Cloyes et al., 2010). Additionally, the episodic nature of traditional therapy fails to provide immediate support during critical moments, significantly limiting its practical effectiveness (Sinha, 2011). Previous studies primarily document these issues through surveys or retrospective interviews (Di Lorenzo et al., 2014; Cloyes et al., 2010), rarely addressing real-time motivational dynamics that influence patient decision-making during recovery. To address this, we explicitly models real-time patient resistance and motivational shifts using dynamic conversational simulations, offering deeper insights into persuasive therapeutic interactions and decision-making processes.
2.2 LLMs for Behavioral Health Applications
Recent advances in LLMs have demonstrated remarkable capabilities in naturalistic human interactions, making them particularly suitable for healthcare communication tasks (Estrellado and Coy, 2025; Laymouna et al., 2024). AI-driven conversational systems, such as Therabot, have successfully enhanced patient engagement and treatment adherence in mental health interventions through personalized, empathetic dialogue (Cai et al., 2023; Gao et al., 2024; Bakoyiannis, 2025). In clinical contexts, LLMs have been employed for mental health triage, diagnostic reasoning, and clinical documentation, providing significant efficiency gains (Park et al., 2024; Fan et al., 2024). However, these applications often overlook critical aspects of long-term behavioral change, particularly in high-risk relapse domains such as addiction treatment. Few existing AI systems explicitly incorporate sustained motivational strategies or therapeutic logic from CBT or MI within longitudinal conversational frameworks (Park et al., 2024; Fan et al., 2024). To fill this gap, our ChatThero framework integrates evidence-based therapeutic methodologies into adaptive, multi-turn conversational agents explicitly designed for long-term addiction recovery support.
2.3 Multi-Agent Simulation Frameworks in Healthcare
Multi-agent simulations have recently emerged as powerful tools for modeling complex healthcare interactions, effectively simulating dynamic clinical scenarios involving multiple stakeholders (e.g., patients, clinicians, environments) (Park et al., 2024; Yu et al., 2024; Karthikesalingam and Natarajan, 2024). Frameworks like AgentClinic and AMIE utilize language agents to replicate clinical decision-making processes, improving diagnostic accuracy and clinical training effectiveness (Park et al., 2024; Fan et al., 2024). Comprehensive AI-driven simulation environments such as AI Hospital and MedSimAI have demonstrated enhanced clinical education outcomes through adaptive training and realistic scenario generation (Yu et al., 2024; Karthikesalingam and Natarajan, 2024; Fan et al., 2024). Despite these advancements, existing frameworks rarely address addiction recovery’s unique challenges, such as high relapse risk, psychological resistance, and environmental triggers. Moreover, prior work typically lacks integration with validated therapeutic strategies (CBT, MI) crucial for addiction treatment (Park et al., 2024; Fan et al., 2024). In response, our ChatThero system advances multi-agent simulation techniques by explicitly embedding therapeutic strategies and modeling negative environmental influences (e.g., social triggers, stressors) to robustly evaluate conversational interventions under realistic, challenging conditions.
2.4 Virtual Patients and Realistic Agent Modeling
Virtual patients (e.g., computer-generated simulations of patient behaviors) have long supported clinical education (for Behavioral Health Statistics and Quality, 2022; Huang et al., 2007; Gordon et al., 2001; Good, 2003; Garrett et al., 2010). Initially rule-based, recent advancements enable LLM-driven virtual patients to simulate complex human behaviors realistically, capturing nuanced psychological and social dynamics (Campillos-Llanos et al., 2021; Lee et al., 2024; Wang et al., 2024; Park et al., 2023). For example, Park et al. demonstrated that LLM-driven virtual agents accurately replicated real individuals’ attitudes and behaviors with high fidelity (Park et al., 2023). In clinical contexts, advanced LLM agents successfully simulate patient-provider interactions, enabling realistic clinical decision-making scenarios and robust agent-driven training platforms (Park et al., 2024; Schmidgall et al., 2024; Fan et al., 2024). Building upon these innovations, our work introduces sophisticated virtual patient agents specifically tailored for addiction treatment, dynamically evolving their profiles and memories to realistically reflect addiction’s psychological complexity. These virtual patients form the core of our multi-agent framework, enabling controlled, ethical, and scalable research into therapeutic dialogue strategies for addiction recovery.
Figure 2: (Left) Each AI patient is initialized with a structured profile including demographic and social context, drug usage history, psychological traits (e.g., impulsivity, distrust), and treatment barriers (e.g., stigma, emotional reliance). These profiles inform role-play-based conversations for the simulated patients. (Middle) Synthetic multi-turn dialogues are generated for each CBT-grounded strategy (e.g., expert endorsement, affirmation), enabling fine-grained persuasion modeling. (Right) ChatThero is optimized through supervised instruction fine-tuning (Stage 1) and DPO-based alignment (Stage 2), where human and AI evaluators provide feedback on response empathy, clarity, and persuasive effectiveness to refine therapeutic behaviors.
@@ 3 Method
ChatThero is an autonomous conversational agent designed to facilitate long-term behavior change and therapeutic support in addiction recovery. Our methodological framework uniquely integrates advanced large language model (LLM)-driven multi-agent simulations, synthetic data generation, clinically-grounded therapeutic strategies, and rigorous evaluation methods (Figure 2). Below, we detail the three core stages of our approach: (1) Patient Profile Generation, (2) Synthetic Data Construction via Multi-Agent Simulations, and (3) Two-Stage Training (SFT + DPO).
@@ 3.1 Patient Profile Generation
Recent studies demonstrated the potential of LLMs to simulate realistic patient behaviors and psychological profiles Yao and Yu (2025); Louie et al. (2024); Du et al. (2024); Yu et al. (2024); Lim et al. (2024). Building upon these advances, our patient simulation consists of a structured Profile and dynamically evolving Memory.
Structured Profile.
@@
Patient profiles are generated via an ethically-informed, multi-stage pipeline based on publicly available narratives from Reddit communities related to addiction (e.g., r/leaves, r/addiction). To ensure privacy, we first prompt a large language model (GPT-4o) to identify and remove any potentially privacy-sensitive or personally identifiable information (PII) from the original posts. After this privacy filtering step, the model summarizes key behavioral indicators and psychological patterns (e.g., sleep disturbances, social isolation) using structured extraction prompts (see Appendix Table 10). Finally, GPT-4o synthesizes a structured profile consisting of:
  • Personality Traits (e.g., impulsivity, emotional reactivity)
  • Substance Use History (e.g., duration, frequency, relapse history)
  • Significant Life Events (e.g., job loss, social isolation)
  • Hypothesized Motivations for Substance Use (e.g., stress relief, coping with insomnia)
This multi-step approach allows the LLM to generate plausible yet synthetic patient profiles that capture behavioral realism while ensuring that no actual Reddit user’s profile is directly reproduced or reused. As a result, the profiles support downstream experimentation while upholding high ethical and privacy standards.
Dynamic Memory.
@@
Each patient agent maintains dynamically evolving memory, recording interactions, emotional states, coping mechanisms, and perceived environmental influences (e.g., relationships, peer pressures, life stressors). An environment agent periodically injects external simulated events (e.g., job loss, relationship breakdown) into the patient’s memory stream, realistically altering motivations and behaviors.
3.2 Synthetic Data Construction via Multi-Agent Simulations
Due to limited real-world addiction treatment conversation data, we developed synthetic dialogues using a structured multi-agent simulation framework (Figure 2, middle panel). Our approach employs patient profiles and therapeutic strategies, generating two complementary datasets:
SFT Dataset.
GPT-4o synthesized multi-turn therapeutic dialogues conditioned directly on patient profiles, employing clinically-validated strategies including MI, CBT, and harm reduction (specific generation prompts in Appendix Tables 5 and 6). Dialogues consist of approximately 50 conversational turns, mirroring realistic therapeutic depth.
DPO Dataset.
Further conversational effectiveness is achieved via multi-agent simulation comprising:
  • Patient Agent (GPT-4o-mini): Dynamically simulates patient resistance and behavioral patterns from structured profiles and memory states (prompt template: Appendix Table 9).
  • Therapy Agent (ChatThero): Deploys adaptive therapeutic strategies from a predefined list of 18 clinically-grounded approaches (Appendix Table 7).
  • Environment Agent: Introduces external stressors, triggering realistic variability in patient responses and conversational contexts.
For each dialogue state, multiple therapeutic responses were generated and ranked by GPT-4o or licensed clinical experts, forming preference pairs for subsequent DPO training (ranking and annotation protocol in Appendix Table 11). This setup allowed controlled yet realistic variations in patient-therapist interactions, significantly enhancing synthetic dataset quality and clinical realism.
3.3 Two-Stage Training (SFT + DPO)
SFT.
Initial conversational capabilities of ChatThero were fine-tuned using the SFT dataset, embedding empathetic communication skills, persuasive strategy knowledge, and realistic dialogue flows derived from structured profiles (generation details in Appendix Tables 56).
DPO.
Using the multi-agent simulation-generated DPO dataset, ChatThero was further refined. Preference pairs guided the model in selecting optimal therapeutic strategies, adaptively adjusting to evolving patient behaviors and external stressors. Human annotations from licensed therapists (300 dialogues, annotation protocol in Appendix Table 13) provided rigorous clinical grounding alongside GPT-4o evaluation.
3.4 Conversation Quality and Clinical Evaluation Framework
4 Experiment

### v2

# ChatThero: A Language Agent for Recovery Support

@@ Abstract
Substance use disorders (SUDs) affect millions of people, and relapses are common, requiring multi-session treatments. Access to care is limited, which contributes to the challenge of recovery support. We present ChatThero, an innovative low-cost, multi-session, stressor-aware, and memory-persistent autonomous language agent designed to facilitate long-term behavior change and therapeutic support in addiction recovery. Unlike existing work that mostly finetuned large language models (LLMs) on patient-therapist conversation data, ChatThero was trained in a multi-agent simulated environment that mirrors real therapy. We created anonymized patient profiles from recovery communities (e.g., Reddit). We classify patients as easy, medium, and difficult, three scales representing their resistance to recovery. We created an external environment by introducing stressors (e.g., social determinants of health) to simulate real-world situations. We dynamically inject clinically-grounded therapeutic strategies (motivational interview and cognitive behavioral therapy). Our evaluation, conducted by both human (blinded clinicians) and LLM-as-Judge, shows that ChatThero is superior in empathy and clinical relevance. We show that stressor simulation improves robustness of ChatThero. Explicit stressors increase relapse-like setbacks, matching real-world patterns. We evaluate ChatThero with behavioral change metrics. On a 1–5 scale, ChatThero raises motivation by +1.71+1.71 points (from $2.39$ to $4.10$) and confidence by +1.67+1.67 points (from $1.52$ to $3.19$), substantially outperforming GPT-5. On difficult patients, ChatThero reaches the success milestone with $26\%$ fewer turns than GPT-5.  [^1] We will release code soon.
@@ 1 Introduction
[…]
Recent work advances theory-grounded counseling dialogues, and there is growing interest in multi-turn evaluation (Lee et al., 2024b; Na, 2024; Qiu et al., 2023). Yet most systems were finetuned on single-session or short context, which does not represent real therapeutic process for recovery. Most evaluations use single-turn quality scores, which miss trajectories and relapse-like setbacks (Li et al., 2025; Zhang et al., 2024b; Hu et al., 2024; Qiu et al., 2023; Sun et al., 2021). In this study, we create a multi-session framework to simulate coherent therapeutic strategies under environmental perturbations and reports cross-session outcomes.
Figure 1: Overview of ChatThero. Left. Multi-stage CBT with MI. The panel shows CBT stages and a parallel MI layer (Spirit; Four Phases—Engage, Focus, Evoke, Plan; OARS micro-skills; core scales and tools). The Therapy Agent picks from a small bank of CBT strategies and MI tools (0–10 scales, E–P–E, decisional balance, values card, agenda mapping) and then combines and reorders them based on the log and current stressors. Stage tags are for orientation purposes, not in a fixed order. The agent infers the current CBT stage and MI phase and sets the next goal. Center. A six-session course shows step-by-step progress. Each session card has a CBT Goal and an MI Focus that matches the phase (e.g., S1: Engage with scaling; S3: Evoke with double-sided reflections; S5: Plan with commitment and barrier planning). Between sessions, an Environment Agent writes stressors, and the Therapy Agent keeps memory and updates the plan. Right. Patient state and profile. Each AI patient starts with a structured Profile (traits, use history, barriers) and a dynamic Memory.
As shown in Figure 1, we frame recovery support as a multi-session, partially observed decision and generation problem with between-session perturbations and craving. We built a simulation based on a realistic environment. (1) A Patient Agent holds a persistent state and resistance level (easy/medium/hard); profiles reflect distributions from recovery communities after careful filtering and anonymization (Valdez & Patterson, 2022). (2) An Environment Agent injects explicit stressors between sessions to shift state (Sinha et al., 2024). (3) A trainable Therapy Agent (ChatThero) chooses and sequences MI/CBT and harm-reduction strategies and keeps memory across sessions (Miller, 1992; McHugh et al., 2010). Training has two stages. First, SFT on clinician-guided synthetic multi-session dialogues learns structure and safety. Second, DPO to decide when to probe, affirm, plan, or switch strategies, with a hard-case curriculum to strengthen carryover under dense stressors.
We evaluate simulated outcomes across sessions. Because MI conceptualizes readiness to change with motivation/confidence rulers and explicitly targets self‐efficacy, motivation and confidence are clinically meaningful proximal outcomes for substance-use treatment (of Health et al., 2019). We therefore measure: (1) motivation and confidence trajectories; (2) Time-to-Success (turns to reach a predefined milestone), with emphasis on high-resistance patients; (3) robustness under explicit stressor injections; and (4) agreement between a rubric-anchored LLM-as-Judge and blinded clinicians. In our study, ChatThero improves motivation from $2.39$ to $4.10$ and confidence from $1.52$ to $3.19$, and it reduces Time-to-Success on hard cases by $26\%$ turns vs. GPT-5, while scoring higher on empathy and clinical relevance.
Our key contributions include: 1) A multi-session, stressor-aware, memory-persistent language-agent formulation with a reproducible simulation. 2) A data–environment construction process: patient profiles from recovery forums and explicit stressor processes that shift state. 3) A two-stage training recipe with a hard-case curriculum for multi-session strategy learning (SFT $\rightarrow$ DPO). 4) An outcomes-focused multi-session evaluation suite, including motivation/confidence trajectories, Time-to-Success, stressor robustness, and human–LLM agreement, with safety procedures.
@@ 2 Related Works
Substance Use Disorders: Treatment Gaps and Behavioral Challenges. Evidence-based treatments (e.g., MAT, CBT) reduce relapse and improve functioning (SAM, 2023; Wooten & Pieper, ; McHugh et al., 2010), yet uptake and retention remain low (e.g., only 22.1% with OUD receive care in a year; <50% remain engaged at 6 months) (Cloyes et al., 2010; SAM, 2023). Barriers include stigma, cost, access, and ambivalence (Sinha, 2011; Cloyes et al., 2010), while episodic therapy often misses high-risk moments (Sinha, 2011). We model evolving resistance and motivation via multi-visit, environment-aware simulations to provide continuity between sessions and test strategy effectiveness under changing conditions.
LLMs and Automated Counseling for Behavioral Health LLMs have recently demonstrated strong naturalistic interaction capabilities, making them suitable for health communication (Estrellado & Coy, 2025; Laymouna et al., 2024). AI conversational systems, such as Therabot (Bakoyiannis, 2025), can enhance engagement and adherence through personalized, empathetic dialogue (Cai et al., 2023; Gao et al., 2024), while in clinical practice, LLMs already support triage, diagnostic reasoning, and documentation with efficiency gains (Park et al., 2024; Fan et al., 2024).
Motivational Interviewing (MI) and Cognitive Behavioral Therapy (CBT). MI Rollnick (2002); Miller & Rose (2009) and CBT Beck (2019) align with Marlatt’s relapse-prevention framework Collier & Marlatt (1995); Frances et al. (2005); Marlatt & Donovan (2005). MI emphasizes empathy, discrepancy, rolling with resistance, and self-efficacy, operationalized via open questions, affirmations, reflections, and eliciting change talk. CBT has strong evidence for substance-use treatment, but in-person delivery and workforce training limit scalability; program requirements can also reduce retention Gertner et al. (2022). Existing chatbots rarely implement MI/CBT explicitly; by contrast, ChatThero encodes MI/CBT tactics and learns strategy selection via reinforcement learning within a simulated environment.
Multi-Agent Simulation, Virtual Patients, and Automated Counseling Multi-agent simulators model patients, clinicians, and environments; systems like AgentClinic, AMIE, AI Hospital, and MedSimAI use such setups for training and education (Park et al., 2024; Schmidgall et al., 2024; Karthikesalingam & Natarajan, 2024; Fan et al., 2024; Yu et al., 2024). Yet most focus on diagnosis or documentation—not relapse-sensitive behavior change—and seldom include explicit stressors, persistent resistance, or validated MI/CBT tactics (Park et al., 2024; Fan et al., 2024). Virtual patients have long supported medical education and now include LLM-based agents capturing social/psychological dynamics (Gordon et al., 2001; Good, 2003; Garrett et al., 2010; Huang et al., 2007; Campillos-Llanos et al., 2021; Park et al., 2023; Lee et al., 2024a; Wang et al., 2024; Park et al., 2024). With real counseling data scarce, recent work turns to synthetic/reconstructed dialogues (PsyQA, SMILE, CBT-LLM, CPsyCoun, HealMe, CACTUS) (Sun et al., 2021; Qiu et al., 2023; Na, 2024; Zhang et al., 2024b; Xiao et al., 2024; Lee et al., 2024a), but typically in single-session, short-context, static-prompt settings (Park et al., 2024; Fan et al., 2024). In contrast, we embed MI/CBT into multi-session simulations, add social triggers and stressors, and build addiction-tailored virtual patients with profiles and evolving memory to test strategy choice, carryover, and robustness.
@@ 3 Method
We model ChatThero as a multi-session, multi-agent system (Figs. 12) with a Patient Agent (persistent state; easy/medium/hard resistance), an Environment Agent injecting between-session stressors, and a trainable Therapy Agent (ChatThero) that selects/sequences MI/CBT/harm-reduction strategies while maintaining longitudinal memory. This setup carries information across visits, adapts plans to adherence signals and stressors, and advances stage-wise goals. The pipeline includes: (1) Patient Profile Generation, (2) Synthetic Data via Multi-Agent Simulation, and (3) Two-Stage Training (SFT $\rightarrow$ DPO).
@@ 3.1 Patient Profile Generation
Recent studies demonstrated the potential of LLMs to simulate realistic patient behaviors and psychological profiles Yao & Yu (2025); Louie et al. (2024); Du et al. (2024); Yu et al. (2024); Lim et al. (2024). Building upon these advances, our patient simulation consists of a structured Profile and dynamically evolving Memory.
@@
Structured Profile. Patient profiles are generated via an ethically-informed, multi-stage pipeline based on publicly available narratives from Reddit communities related to addiction (e.g., r/leaves, r/addiction). To ensure privacy, we first prompt a local LLM to identify and remove any potentially privacy-sensitive or personally identifiable information (PII) from the original posts. After this privacy filtering step, the model summarizes key behavioral indicators and psychological patterns (e.g., sleep disturbances, social isolation) using structured extraction prompts (see Table 9). Finally, GPT-4o synthesizes a structured profile consisting of: 1) Personality Traits (e.g., impulsivity, emotional reactivity); 2) Substance Use History (e.g., duration, frequency, relapse history); 3) Significant Life Events (e.g., job loss, social isolation); 4) Hypothesized Motivations for Substance Use (e.g., stress relief, coping with insomnia). This multi-step approach allows the LLM to generate plausible yet synthetic patient profiles that capture behavioral realism while ensuring that no actual Reddit user’s profile is directly reproduced or reused. As a result, the profiles support downstream experimentation while upholding high ethical and privacy standards.
@@
Dynamic Memory. Each patient agent maintains a dynamically evolving memory, recording interactions, emotional states, coping mechanisms, and perceived environmental influences (e.g., relationships, peer pressures, life stressors). An environment agent periodically injects external simulated events (e.g., job loss, relationship breakdown) into the patient’s memory stream, realistically altering motivations and behaviors.
3.2 Single-Session Synthetic Data (Agents & Objectives)
Because real-world addiction-treatment dialogues are scarce and sensitive, we first construct single-session synthetic data to teach core therapeutic skills within one encounter (Figure 2, middle).
Figure 2: Data synthesis and two-stage training for ChatThero. Left: A Therapy Agent and a Patient Agent generate multi-session dialogues while an Environment Agent injects stressors, prompting strategy switches. Right: Stage-1 SFT teaches safe MI/CBT structure; Stage-2 DPO refines timing and strategy selection. At test time, the agent keeps memory across sessions, adapts after stressors, and targets higher motivation/confidence with lower time-to-success.
Agents and MI/CBT usage (single-session). We instantiate two agents: (i) a Patient Agent (PA) that presents substance-use–specific concerns, goals, and resistance from a structured profile (motivation, confidence/self-efficacy, current craving level, high-risk triggers/contexts, adherence to plans or MOUD, recent lapses) and an explicit difficulty label {easy, medium, hard}. Here, easy denotes patients who are willing to accept treatment and actively cooperate; medium denotes patients who show ambivalence with noticeable resistance to guidance; and hard denotes long-term users who distrust clinicians and struggle to resist drug-use urges. The PA surfaces these states each turn as a compact summary to guide the therapist policy. (ii) A trainable Therapy Agent(TA) that conducts an MI/CBT-consistent session by selecting from predefined strategy primitives (Appendix Table 6) mapped to MI and/or CBT with explicit schemas. Concretely for MI, the policy targets substance-use change by eliciting change talk around personally valued goals (health, legal, family), running decisional balance on use vs. not-use days, using confidence/importance rulers for quit/cut-down plans, rolling with resistance about abstinence vs. harm-reduction, and shaping commitment language for next-step actions. For CBT, the policy operates on drug-use episodes via functional analysis (antecedent–behavior–consequence), trigger mapping (people/places/things), cognitive reappraisal of urges (e.g., “I can’t sleep without using”), coping-skill rehearsal for cravings (delay, distract, de-stress, grounding), refusal-skills scripting, stimulus control (avoiding procurement cues), graded exposure to high-risk cues with urge-surfing, and problem-solving for access/adherence barriers (appointments, transportation).
Interaction protocol. Each dialogue follows a substance-use–specific agenda within a fixed-turn budget ($\sim$60 turns): rapport & goal alignment (values, quit/cut-down intent, MOUD preferences) $\rightarrow$ episode clarification (concise functional analysis of a recent use/near-use: antecedents, triggers, craving peak, consequences) $\rightarrow$ plan formulation (translate insights into one or two actionable steps) $\rightarrow$ next-step micro-commitment. In lieu of generic “homework,” the session ends with a concrete, time-bounded micro-assignment tied to risk and feasibility—e.g., a trigger/urge log (time, place, people, intensity), a three-line coping card for an anticipated window, a single refusal-line rehearsal with a specific peer, a stimulus-control action (remove a procurement contact/app). The Patient Agent provides a compact state summary each turn (craving, trigger salience, confidence, recent lapse flags), and the Therapy Agent generates the utterance plus a measurable micro-commitment (deadline and success criterion).
Safety and fidelity controls. All turns pass rule-based filters (diagnosis prohibition, medication disclaimers, crisis routing) and guardrails to penalize off-policy patterns (e.g., argumentation, prescriptive advice without permission).
3.3 Multi-Session Synthesis
To address single-session limitations, we extend synthesis to multi-session episodes (3–6 outpatient visits) and explicitly stage therapy across sessions.
Therapy staging across sessions (strategy evolves). The TA does not use a static policy; it follows the six CBT stages shown in Fig. 1 (Left) and operationalized in Fig. 1 (Center): (S1) Build trust & assessment — set safety/limits and begin a one-page trigger/urge log; MI focus: engage and use 0–10 rulers for importance/confidence. (S2) Identifying negative cognitions — review the log and map thought $\rightarrow$ feeling $\rightarrow$ body $\rightarrow$ action; MI focus: values/agenda mapping to sharpen goals. (S3) Challenging false beliefs — name the core thought (“one time won’t hurt”), draw an urge curve, and reality-test; MI focus: evoke change talk (DARN) with double-sided reflections. (S4) Restructuring cognitive patterns — create an If–Then plan with grounding and partner support; MI focus: bridge Evoke $\rightarrow$ Plan (CAT: commitment, activation, taking steps). (S5) Behavioral skill building — redesign a high-risk window (e.g., drawing + paced breathing) with a crisis micro-plan and backup steps; MI focus: planning with barrier-anticipation. (S6) Consolidation & termination — review gains, extend the plan (four more weeks), add relapse-prevention and peer support; MI focus: summarize accumulated change talk and set maintenance. Across sessions, the PA maintains a persistent memory $m_{t}$ (craving, trigger salience, motivation, self-efficacy, adherence, lapse flag), summarized to TA at the start of each session so that strategy choice advances the current stage rather than treating sessions as isolated events.
Environment Agent and rule-based memory updates (Fig. 2). Between sessions, an Environment Agent(EA) injects stressors from a curated catalog (short NL description, severity=low/medium/high, duration=days–weeks), organized as: (a) peer/availability (invitations, proximity to using peers, easier access), (b) work/academic (deadlines, shift changes, supervisor pressure), and (c) home/context (family conflict, housing instability, stigma). Each event is recorded in a stressor ledger $\mathcal{L}_{t}$ and applied to the persistent patient memory $m_{t}$ via interpretable, rule-based updates—e.g., peer exposure $\uparrow$ trigger salience; sustained workload $\downarrow$ self-efficacy; supportive events or successful refusals $\uparrow$ motivation/confidence; thresholded accumulation toggles lapse_flag. At the next visit, TA receives the $(m_{t},\mathcal{L}_{t})$ summary, references prior stage artifacts (log, thought record, coping card, plan), and adjusts the staged goal (S1–S6) accordingly—mirroring the “mental state after session” trajectory in Fig. 1 (Right). The whole detailed clinical scenarios settings are shown in Appendix A.2
3.4 Training (SFT & DPO): Objectives, Data Construction, and Procedure
Supervised Fine-Tuning (SFT). We first supervise ChatThero on single-session dialogues synthesized by GPT-4o from structured patient profiles, using prompts that require MI/CBT/harm-reduction grounding and safety boundaries (Appendix Tables 45); each profile yields multiple scenarios for lexical/situational diversity. The target is the therapist reply $y$ given user turn $x$ and the PA state summary $s$, including an internal strategy tag/rationale (hidden at inference). We then extend SFT to multi-session episodes (3–6 sessions) that explicitly reference prior plans/outcomes and update state under between-visit stressors; inputs are augmented to $(x,s,m_{t},\mathcal{L}_{t})$ so the model learns cross-visit carryover (artifact updates, plan continuation). Training uses standard token-level cross-entropy, mixing single-session and multi-session examples in one curriculum.
Direct Preference Optimization (DPO). For single-session preference learning, a lightweight simulation pairs a PA (GPT-4o-mini) expressing profile-aligned resistance with candidate TA replies drawn from 18 predefined MI/CBT strategy primitives (Appendix Table 6). Candidates are ranked by GPT-4o using a rubric (empathy, clinical relevance, safety, concreteness), producing preference pairs (Appendix Table 10); we optimize a standard DPO objective with the SFT policy as reference. For multi-session DPO, we form longitudinal preference pairs at cross-visit decision points (relapse framing, strategy switching, plan revision under new stressors, correct use of prior information) and expand the rubric with carryover fidelity, stressor alignment, and stage coherence (Appendix Table 12); these items are up-weighted when forming preferences. This two-phase (single-session $\rightarrow$ multi-session) DPO improves within-visit quality and multi-session persistence without requiring an explicit reward model.  [^2] We filter generations with safety/ethics rules, retain only privacy-preserving synthesized profiles, and de-duplicate near-duplicates. Hyperparameters and sampling/ranking templates are detailed in Appendix 812.
4 Experimental Setup